# LLM-3D Print: Large Language Models To Monitor and Control 3D Printing

### v2

@@ 2.3 Planning Agent
While LLMs have demonstrated emergent reasoning capabilities, their performance can be significantly enhanced when guided by a structured reasoning framework 102, 103. Additionally, because the output of an LLM is heavily influenced by the quality and specificity of the prompts, it becomes essential to adapt these prompts based on the specific problems being addressed. Providing a well-defined reasoning structure and carefully tailored prompts allows the LLM to deliver more accurate and effective solutions, ensuring better outcomes for complex tasks104.
To accomplish this, the planning agent (Fig 3) (b) is structured with two internal modules. The first module analyzes known information, observations and detected failures to select, adapt, and refine standard reasoning prompts (SI 2). This ensures that the planner works with the most optimized and relevant prompts when formulating a plan. The second module then leverages these tailored prompts and reasoning frameworks to develop a specific, actionable plan. This plan is carried out by the executor agents, which interact with the 3D printer via API to gather necessary data and implement the solution. This dual-module design enables the planning agent to effectively and efficiently resolve any issues that arise during
@@ 2.4 Executor Agent
The Agent Executor (Fig. 3c) , a crucial component of the framework, utilizes the ReAct (Reasoning and Acting) method 105 to execute the generated plans.
@@ 3 Experimental Setup
The printer setup consists of a consumer grade 3D printer modified to stream in-situ images and system information while also exposing an Application Program Interface (API) to accept commands to dynamically adjust printing parameters under the direction of an LLM agent. This is achieved with Klipper, an open source firmware project that delegates the task of parsing G-code and process monitoring away from the controller board to an external computer. Klipper in conjunction with ancillary plugins such as Crowsnest, Moonraker, and Mainsail provide a cohesive interface which allows for monitoring and control over various aspects of the printing process.
A stock Creality Ender 5 Plus 3D printer was reflashed with Klipper and configured alongside a standard desktop computer used as the host machine to deliver commands to the printer’s motor control unit. Two SVPRO1080P cameras with a 2.8 mm to 12 mm range of manual focus were placed to provide top and front views of the print (Fig. 2b). These cameras are configured with the Crowsnext plugin to stream images of the printing process from the host computer. Of the two cameras, the top one captures in-situ images of the printing process and these images are then provided as input to the multimodal LLM to identify and evaluate any defects in the ongoing print. An input image is captured using with this camera at the specified end of a sequence of GCode commands. During the capture process the print is paused and toolhead is parked at home positions to avoid toolhead interference over the printed segment. These process parameter changes are made using the Moonraker API which exposes access to configurations such as the toolhead temperature, print speed, extrusion rate, or fan speed to name a few.
@@ 4.1 Multi Layer Prints
Additionally, layer-wise analysis in Figures 4(c and g) reveals continuous improvement in layer quality and infill patterns in the LLM-optimized prints, whereas the baseline prints (Figures 4(d and h)) deteriorate as the print progresses. The LLM-optimized approach not only addresses issues such as rough edges, inconsistent surfaces, and poor layer adhesion but also significantly enhances overall print quality, resulting in superior final products.
@@ 4.1.1 Parameter Optimization
This metric provides a measure of how well the printed layer covers the intended area. For single-layer prints, the LLM identified several key parameters to adjust, including print speed, flow rate, pressure advance, and retraction flow for PLA prints, and nozzle temperature for TPU prints. Fig 5 illustrate the changes made by the LLM for single layer, with Figure 5a PLA print and Figure 5b TPU print. For print speed, the LLM suggested reducing it to 75% to improve filament adhesion and minimize gaps. This change led to smoother deposition and more consistent layers. An adjustment to the flow rate, increasing it slightly above 100%, addressed the under-extrusion issues observed in the initial segments, resulting in a more uniform filament flow. A minor increase in pressure advance (up to 0.1) helped in managing the flow of filament, especially in corners and intricate details, reducing stringing and blobs. Optimizing retraction flow to around 25% mitigated the issues related to oozing and stringing between travel moves, enhancing the print’s clean lines and precision. For TPU, the LLM suggested raising the nozzle temperature to 220°C, which improved filament melting and flow, leading to better layer adhesion and reduced stringing. The continuous increase in occupancy throughout the single-layer print, as shown, demonstrates the effectiveness of the LLM’s optimization strategies in improving print quality.
@@
For multi-layer print of wrench as illustrated in Figure 5 c, the LLM accurately identifies that the print speed and extrusion rate are off-nominal and actively seeks to optimize these parameters throughout the printing process. Additionally, despite the printer bed being correctly calibrated, the LLM continuously monitors the Z-axis position. Based on the provided layer height and the current layer number, the LLM adjusts the Z-offset of the printer to ensure a smooth surface finish. Notably, at Layer 10, the LLM detected that the Z position was misaligned and persistently attempted to correct it, demonstrating its capability to adapt and fine-tune the printing parameters.
For the print of raised text Figure 5d, the LLM initially identifies signs of under-extrusion and responds by increasing the flow rate, adjusting extruder acceleration, and reducing the print speed. These initial adjustments aim to stabilize the print quality and address the immediate under-extrusion issue. After these corrections, the LLM enters an optimization phase, where it systematically fine-tunes the print speed and flow rate to find the optimal settings for the specific print conditions.
@@ 5.1 Analysis of LLM-generated Response
As illustrated in Fig.6, the expert-annotated figure identifies four key failure modes: (1) extrusion inconsistency, (2) stringing, (3) layer separation, and (4) a material blob. The LLM successfully identified the first three issues and proposed actionable corrections. However, it did not flag the blob artifact, suggesting some limitations in recognizing less distinct defect types. The blob, formed by excess filament ending in fine strands, visually resembles stringing, which may have led to its misclassification. The model’s diagnostic reasoning was guided by domain-relevant questions generated through LLM-refined structured prompting (e.g., ”What could be causing stringing and poor adhesion?”) and did not rely on retraining or fine-tuning.
@@ 5.2 LLM Identified Defects
Error detection is a vital component of the proposed closed-loop framework for additive manufacturing. While LLMs are highly effective at processing textual data and sensor logs, detecting visual or geometric print defects remains a significant challenge, particularly in the absence of fine-tuning on 3D printing-specific datasets. To assess this capability, we evaluated the LLM’s performance in identifying print defects during the fabrication of a single layer, comparing its outputs with expert annotations (SI: 8). In the absence of standardized metrics for print quality assessment, such comparisons provide a practical basis for evaluation. Prior work has employed annotated datasets and precision-recall analysis to assess model performance in similar contexts 38, 106, and we adopt a comparable approach.
@@
Figure 7b presents a comparative analysis of participant responses, expert annotations, and the LLM’s detection results. The analysis shows that both the LLM and the majority of participants, including experts, consistently identified stringing and extrusion-issues as prominent failure modes across multiple layers. In addition, the LLM was able to detect more complex defects, such as bed adhesion failures, warping, and layer separation, which were sometimes overlooked by less experienced participants. These results suggest that the LLM not only matches near-human-level diagnostic capabilities in common error scenarios but can also offer enhanced detection in more subtle or cumulative failure conditions, reinforcing its role as a reliable agent in automated quality monitoring.
@@
A more quantitative comparison of detection performance is provided in Figure 8, which presents confusion matrices for both the LLM and human survey responses against expert annotations. The LLM demonstrates strong precision and recall in identifying high-impact defects such as extrusion issues and stringing/oozing, with few false positives or false negatives. In contrast, human survey responses showed greater variability, particularly in the classification of subtle defects such as ghosting, elephant foot, and print cracks. Participants frequently over-identified these rare defects, leading to a higher rate of false positives. This further highlights the LLM’s ability to maintain diagnostic accuracy while minimizing over-attribution.
@@ 5.3 Generalization Without Supervision
Recent efforts in automated defect detection have primarily focused on training deep learning models with large-scale annotated datasets, yet these efforts often optimize only for a narrow range of simple failure modes. For instance, one approach trained a convolutional neural network (CNN) on over 120,000 labeled images, using only three coarse defect categories: “Good-quality,” “Under-extrusion,” and “Over-extrusion” 38. Similarly, a vision transformer (ViT)-based model relied on heavily curated data and extensive training to identify a limited set of extrusion defects 106. While such models perform well within their labeled domains, their effectiveness is bounded by the specificity of the annotations. They struggle to generalize to new failure types, printer hardware, or parameter variations not explicitly present in the training data. While the LLM may occasionally misattribute certain defects on a per-layer basis, such as confusing blobs with stringing or overgeneralizing failure categories, its aggregate performance across entire prints leads to significantly improved structural outcomes. By consistently identifying key defect patterns and applying corrective parameter adjustments, the LLM contributes to enhanced print quality over time.
@@
Our framework was deployed on a standard consumer-grade FDM printer with a 1 mm nozzle, lacking manufacturer-provided settings. The LLM successfully diagnosed and corrected print issues without device-specific tuning. Its reasoning capabilities extended across incomplete inputs and generalized settings. The framework also adapts dynamically to available sensor data, including temperature, retraction settings, print speed, and G-code offsets, based on firmware accessibility. This sensor-agnostic design ensures compatibility with diverse printer models and configurations, supporting wide applicability without requiring hardware-specific integration.

### v3

@@ 2.3 Planning Agent
While LLMs have demonstrated emergent reasoning capabilities, their performance can be significantly enhanced when guided by a structured reasoning framework 106, 107. Additionally, because the output of an LLM is heavily influenced by the quality and specificity of the prompts, it becomes essential to adapt these prompts based on the specific problems being addressed. Providing a well-defined reasoning structure and carefully tailored prompts allows the LLM to deliver more accurate and effective solutions, ensuring better outcomes for complex tasks108.
To achieve this, the planning agent (Fig. 3b) is organized into two modules. The adapter module analyzes observations and detected failures to adjust standard reasoning prompts (SI 2), tailoring them to the specific context while refining and optimizing their relevance. The planner module then leverages these adapted prompts and reasoning frameworks to generate a concrete, actionable plan. Executor agents carry out this plan by interacting with the 3D printer via API to collect data and implement solutions. This dual-module design enables the planning agent to address issues effectively and efficiently during operation.
@@ 2.4 Executor Agent
The Agent Executor (Fig. 3c) , a crucial component of the framework, utilizes the ReAct (Reasoning and Acting) method 109 to execute the generated plans.
@@ 3 Experimental Setup
The printer setup consists of a consumer grade 3D printer modified to stream in-situ images and system information while also exposing an Application Program Interface (API) to accept commands to dynamically adjust printing parameters under the direction of an LLM agent. This is achieved with Klipper19, an open source firmware project that delegates the task of parsing G-code and process monitoring away from the controller board to an external computer. Klipper 19 in conjunction with ancillary plugins such as Crowsnest110, Moonraker111, and Mainsail112 provide a cohesive interface which allows for monitoring and control over various aspects of the printing process.
A stock Creality Ender 5 Plus 113 and Creality Ender 3 114 3D printer was reflashed with Klipper (v0.13.0)19 and configured alongside a standard desktop computer used as the host machine to deliver commands to the printer’s motor control unit. Two SVPRO1080P cameras with a 2.8 mm to 12 mm range of manual focus were placed to provide top and front views of the print (Fig. 2b). These cameras are configured with the Crowsnest (v4.1.16)110 plugin to stream images of the printing process from the host computer. Of the two cameras, the top one captures in-situ images of the printing process and these images are then provided as input to the multimodal LLM to identify and evaluate any defects in the ongoing print. An input image is captured using with this camera at the specified end of a sequence of GCode commands. During the capture process the print is paused and toolhead is parked at home positions to avoid toolhead interference over the printed segment. These process parameter changes are made using the Moonraker (v0.9.3)111 API which exposes access to configurations such as the toolhead temperature, print speed, extrusion rate, or fan speed, to name a few.
Our multi-agent hierarchical framework is implemented in Python (v3.12.4) and builds upon LangChain (v0.3.27) 115 along with LangGraph (v0.6.7) 116. To demonstrate that our framework is robust to shifts in model behavior over time 117, we evaluated it using two large language models: ChatGPT-4o 118 and ChatGPT-4.1 119. This ensures that our results are not tied to one specific model version, but generalize across newer and evolving LLMs.
@@ 4.1 Multi Layer Prints
Additionally, layer-wise analysis in Fig. 4(c and g) reveals continuous improvement in layer quality and infill patterns in the LLM-optimized prints, whereas the baseline prints (Figures 4(d and h)) deteriorate as the print progresses. The LLM-optimized approach not only addresses issues such as rough edges, inconsistent surfaces, and poor layer adhesion but also significantly enhances overall print quality, resulting in superior final products.
@@ 4.2 Parameter Optimization
This metric provides a measure of how well the printed layer covers the intended area. For single-layer prints, the LLM identified several key parameters to adjust, including print speed, flow rate, pressure advance, and retraction flow for PLA prints, and nozzle temperature for TPU prints. Fig. 5 illustrate the changes made by the LLM for single layer, with Fig. 5a PLA print and Fig. 5b TPU print. For print speed, the LLM suggested reducing it to 75% to improve filament adhesion and minimize gaps. This change led to smoother deposition and more consistent layers. An adjustment to the flow rate, increasing it slightly above 100%, addressed the under-extrusion issues observed in the initial segments, resulting in a more uniform filament flow. A minor increase in pressure advance (up to 0.1) helped in managing the flow of filament, especially in corners and intricate details, reducing stringing and blobs. Optimizing retraction flow to around 25% mitigated the issues related to oozing and stringing between travel moves, enhancing the print’s clean lines and precision. For TPU, the LLM suggested raising the nozzle temperature to 220°C, which improved filament melting and flow, leading to better layer adhesion and reduced stringing. The continuous increase in occupancy throughout the single-layer print, as shown, demonstrates the effectiveness of the LLM’s optimization strategies in improving print quality.
@@
For multi-layer print of wrench as illustrated in Fig. 5 c, the LLM accurately identifies that the print speed and extrusion rate are off-nominal and actively seeks to optimize these parameters throughout the printing process. Additionally, despite the printer bed being correctly calibrated, the LLM continuously monitors the Z-axis position. Based on the provided layer height and the current layer number, the LLM adjusts the Z-offset of the printer to ensure a smooth surface finish. Notably, at Layer 10, the LLM detected that the Z position was misaligned and persistently attempted to correct it, demonstrating its capability to adapt and fine-tune the printing parameters.
For the print of raised text Fig. 5d, the LLM initially identifies signs of under-extrusion and responds by increasing the flow rate, adjusting extruder acceleration, and reducing the print speed. These initial adjustments aim to stabilize the print quality and address the immediate under-extrusion issue. After these corrections, the LLM enters an optimization phase, where it systematically fine-tunes the print speed and flow rate to find the optimal settings for the specific print conditions.
@@
4.3 Cross-Platform Validation
Figure 6: Baseline versus LLM-optimized prints. Baseline samples exhibit visible defects, including seam gaps and incomplete bonding at the infill–perimeter interface, whereas LLM-optimized prints display fully consolidated features across (a) square, (b) hexagonal, (c) hemispherical, and (d) auxetic geometries.
To demonstrate the generalizability of our approach across different hardware platforms, we deployed the framework on a Creality Ender 3 printer114, which features a distinctly different sensor configurations compared to the previously tested Ender 5 Plus113. Notably, the Ender 3 lacks the BLTouch automatic bed leveling sensor present in the Ender 5 Plus setup. To accommodate this hardware difference, we simply removed the BLTouch-related parameters and queries from the printer documentation accessible to the LLM agents, demonstrating the framework’s adaptability to varying hardware configurations. This printer setup utilized the same 1mm nozzle diameter; however, since manufacturer default parameters are typically unavailable for non-standard nozzle sizes, we selected print settings that represented reasonable starting points within the operational envelope, though these settings naturally introduced certain processing challenges that would allow the framework to demonstrate its autonomous error detection and correction capabilities. Additionally, to evaluate the framework’s robustness in addressing defects originating from pre-processing stages, we introduced variations in the infill overlap percentage during the slicing phase, a parameter that becomes embedded within the G-code tool path and remains immutable during runtime execution. This scenario reflects common industrial workflows where G-code files may be sourced from external suppliers or generated using different slicer configurations for intellectual property protection, as manufacturers often receive only the final toolpath data rather than the original design files or slicer parameters. Such situations can result in suboptimal layer adhesion characteristics that cannot be directly corrected through runtime parameter adjustments, thereby testing the framework’s ability to compensate for fixed toolpath limitations through optimization of available dynamic parameters such as extrusion rate, temperature, and printing velocities. Furthermore, to validate that the framework operates independently of visual characteristics such as filament color, we conducted additional prints using PLA material in a different color (RED) from the previous experiments. This color variation test ensures that the image-based defect detection and quality assessment modules function robustly across different visual presentations of the printed material, eliminating potential bias related to specific color properties or contrast characteristics.
Cross-platform validation shows distinct defect patterns in baseline prints, absent in the LLM optimized versions, highlighting the framework’s adaptive correction capabilities. Most prominently, as illustrated in Fig.6, the baseline prints exhibit much better print quality; however, they show consistent gaps at seam positions (the locations where each layer’s perimeter and infill path begins and ends), creating potential weak points due to inconsistent material deposition during nozzle acceleration and deceleration. Additionally, noticeable gaps appear at the transition zones where infill patterns connect to perimeter walls, indicating suboptimal overlap parameters embedded in the G-code during slicing. Significantly, the LLM-optimized prints initially display these identical defects in the first few layers, demonstrating that the framework begins with the same compromised conditions. However, within approximately 5 iterations, the system successfully identifies these manifestations as under-extrusion issues and implements corrective measures by adjusting extruder speed, temperature, and extrusion multiplier parameters to compensate for the fixed toolpath limitations. Unlike the previous experiments that began with deliberately poor parameter selections, this validation used near-default printing parameters, resulting in baseline prints with relatively minimal stringing and fewer catastrophic defects, yet the framework still achieved measurable improvements in seam quality, infill-perimeter bonding, and overall surface consistency. This demonstrates the system’s sensitivity to subtle quality variations and its ability to optimize even moderately acceptable print conditions toward higher precision manufacturing standards.
4.4 Experimental Validation
Figure 7: Compression performance of baseline and LLM-optimized samples. Baseline prints show premature failure with lower peak loads, while LLM-optimized counterparts exhibit enhanced structural integrity and significantly higher peak load capacities: (a) square structure (×5.06), (b) hexagonal structure (×1.60), (c) hemispherical structure (×1.34), and (d) auxetic structure (×2.46).
The observed defects, particularly those occurring at seam positions and infill-perimeter transition zones, represent critical structural vulnerabilities that extend beyond aesthetic concerns. Seam discontinuities create stress concentration points that can initiate crack propagation under mechanical loading, while insufficient infill overlap reduces the effective load transfer between perimeter walls and internal structures. These localized weaknesses can significantly compromise the overall structural integrity of printed components, particularly in applications requiring reliable mechanical performance. To quantify the mechanical implications of these print quality improvements, we conducted compression testing using an Instron universal testing machine equipped with a 100 kN load cell. Four distinct geometries were evaluated: square structure, hemispherical strucutre, auxetic structure, and a hexagonal unit cell — each representing fundamental building blocks commonly employed in structural applications ranging from packaging to metamaterial design. Given the substantial load capacity of the testing apparatus, the sensitivity limitations for detecting minor load variations necessitated a modification to the LLM optimization strategy. The framework prompts were enhanced to include specific mechanical directives: ”strengthen the print in x/y direction for compression test, z is the build direction,” ensuring that parameter optimization prioritized structural performance in addition to visual quality. The compression test results demonstrate substantial mechanical performance improvements across all tested geometries. The square geometry showed the most dramatic enhancement with a 5.06× increase in peak load capacity, rising from approximately 0.8 kN in baseline samples to over 4.0 kN in LLM-optimized prints. The hexagonal unit cell achieved a 1.60× improvement, while the hemisphere and auxetic structures exhibited 1.34× and 2.46× increases, respectively. The load-displacement curves reveal that LLM-optimized samples not only achieve higher peak loads but also demonstrate more consistent mechanical behavior with smoother load progression and more predictable failure patterns. Notably, the baseline samples across all geometries exhibit premature failure and erratic load-displacement relationships, indicating structural inconsistencies that correlate directly with the visual defects identified during the printing process. These results validate that the framework’s visual quality improvements translate directly into enhanced mechanical performance, establishing a quantitative link between defect correction and functional part integrity.
@@ 5.1 Analysis of LLM-generated Response
As illustrated in Fig. 8, the expert-annotated illustration identifies four key failure modes: (1) extrusion inconsistency, (2) stringing, (3) layer separation, and (4) a material blob. The LLM successfully identified the first three issues and proposed actionable corrections. However, it did not flag the blob artifact, suggesting some limitations in recognizing less distinct defect types. The blob, formed by excess filament ending in fine strands, visually resembles stringing, which may have led to its misclassification. Importantly, all diagnostic reasoning was performed in-context using a base LLM (ChatGPT-4o) without any retraining or fine-tuning. The reasoning process was guided by domain-specific, generalized structured prompts (e.g., “What could be the reasoning for failure?”, ”How would you solve the problems without human intervention?”) that we developed in-house.
@@ 5.2 LLM Identified Defects
Error detection is a vital component of the proposed closed-loop framework for additive manufacturing. While LLMs are highly effective at processing textual data and sensor logs, detecting visual or geometric print defects remains a significant challenge, particularly in the absence of fine-tuning on 3D printing-specific datasets. To assess this capability, we evaluated the LLM’s performance in identifying print defects during the fabrication of a single layer, comparing its outputs with expert annotations (SI:4). In the absence of standardized metrics for print quality assessment, such comparisons provide a practical basis for evaluation. Prior work has employed annotated datasets and precision-recall analysis to assess model performance in similar contexts 38, 120, and we adopt a comparable approach.
@@
Fig.9b presents a comparative analysis of participant responses, expert annotations, and the LLM’s detection results. The analysis shows that both the LLM and the majority of participants, including experts, consistently identified stringing and extrusion-issues as prominent failure modes across multiple layers. In addition, the LLM was able to detect more complex defects, such as bed adhesion failures, warping, and layer separation, which were sometimes overlooked by less experienced participants. These results suggest that the LLM not only matches near-human-level diagnostic capabilities in common error scenarios but can also offer enhanced detection in more subtle or cumulative failure conditions, reinforcing its role as a reliable agent in automated quality monitoring.
@@
A more quantitative comparison of detection performance is provided in Fig.10, which presents confusion matrices for both the LLM and human survey responses against expert annotations. The LLM demonstrates strong precision and recall in identifying high-impact defects such as extrusion issues and stringing/oozing, with few false positives or false negatives. In contrast, human survey responses showed greater variability, particularly in the classification of subtle defects such as ghosting, elephant foot, and print cracks. Participants frequently over-identified these rare defects, leading to a higher rate of false positives. This further highlights the LLM’s ability to maintain diagnostic accuracy while minimizing over-attribution.
@@ 5.3 Generalization Without Supervision
Recent efforts in automated defect detection have primarily focused on training deep learning models with large-scale annotated datasets, yet these efforts often optimize only for a narrow range of simple failure modes. For instance, one approach trained a convolutional neural network (CNN) on over 120,000 labeled images, using only three coarse defect categories: “Good-quality,” “Under-extrusion,” and “Over-extrusion” 38. Similarly, a vision transformer (ViT)-based model relied on heavily curated data and extensive training to identify a limited set of extrusion defects 120. While such models perform well within their labeled domains, their effectiveness is bounded by the specificity of the annotations. They struggle to generalize to new failure types, printer hardware, or parameter variations not explicitly present in the training data. While the LLM may occasionally misattribute certain defects on a per-layer basis, such as confusing blobs with stringing or overgeneralizing failure categories, its aggregate performance across entire prints leads to significantly improved structural outcomes. By consistently identifying key defect patterns and applying corrective parameter adjustments, the LLM contributes to enhanced print quality over time.
@@
Our framework was deployed on a standard consumer-grade MEX printer with a 1 mm nozzle, lacking manufacturer-provided settings. The LLM successfully diagnosed and corrected print issues without device-specific tuning. Its reasoning capabilities extended across incomplete inputs and generalized settings. The framework also adapts dynamically to available sensor data, including temperature, retraction settings, print speed, and G-code offsets, based on firmware accessibility. This sensor-agnostic design ensures compatibility with diverse printer models and configurations, supporting wide applicability without requiring hardware-specific integration.
5.4 Data Security
Our architecture strictly enforces a separation of roles: the system can only modify runtime parameters during printing, never altering or overwriting the slicer-generated G-code. This safeguard prevents inadvertent or malicious modifications of toolpaths, which past studies have shown can introduce subtle defects or structural compromises. For example, attackers have demonstrated that small edits to G-code may degrade mechanical performance in additive manufacturing121, 122.
We adopt a modular deployment approach that assigns modules based on their sensitivity and trust requirements. Modules that perform vision-based defect detection and real-time control may be hosted locally using fine-tuned or in-house models; these operate purely on image data and machine state, which contain no proprietary design geometry. General reasoning tasks, such as prompt adaptation, decision planning, and high-level correction logic, can run on external LLM services. Because these reasoning modules consume only abstracted defect descriptors and machine metrics, not original design files or toolpaths, they pose minimal risk of leaking sensitive data.
This partitioning mirrors best practices in secure AI architectures and edge–cloud systems. It allows computation to shift toward external servers for tasks requiring large model capacity, while retaining sensitive processing on-premises to comply with security policies. In highly constrained environments, the framework supports full local deployment so that no external model ever sees proprietary or sensitive data. This flexibility ensures that organizations can balance performance, resource constraints, and confidentiality needs.